OpenStudy (anonymous):

he purchasing director of a bank has just received a shipment of 100 printers. She has been told that the shipment includes 5 defective printers. She randomly selects two printers and tests them. If both printers work the purchasing director accepts the shipment. what is the probability the shipment is rejected?

OpenStudy (ragingsquirrel):

I think it would be 5% divided by two, but dont tae my word for it unless you have to.

OpenStudy (anonymous):

no that doesnt come out to the right answer.... the answer is 0.098 i just dont know how to get to that answer

OpenStudy (anonymous):

try \[\frac{95}{100}\times \frac{94}{99}\] as the probability that neither are defective

OpenStudy (anonymous):

and so \[1-\frac{95}{100}\times \frac{94}{99}\] as the probability that the shipment is accepted

OpenStudy (kropot72):

This can also be solved by finding the hypergeometric probabilities of 1 defective (P(1)=0.09595) and 2 defective (P(2)=0.00202) and adding P(1) + P(2).

OpenStudy (kropot72):

@satellite73 Your comment after your final expression should read "...........as the probability that the shipment is rejected "